OpenStudy (anonymous):

what is 2 4ths and 5 6ths addes together

OpenStudy (anonymous):

2/4 + 5/6 = 1 1/3

OpenStudy (anonymous):

1 and 1/3

OpenStudy (anonymous):

As a decimal, 1.033333333333333 As a fraction 1 1/3 As a percent 1.03%

OpenStudy (ybarrap):

2/4 + 5/6 = (2*6 + 5*4)/4*6 =(12+20)/24 =32/24 =(4*8)/(3*8) =4/3

OpenStudy (anonymous):

Try to get a common denominator. Multiply each equation by a fraction equal to one that would get you the same denominator. In this case, it would be 6/6 and 4/4. 2/4 * 6/6 = 12/24. 5/6 * 4/4 = 20/24. Add 12/24 and 20/24 to get 32/24. Simplify to get 4/3, or 1 1/3, or 1.33333 etc.

OpenStudy (anonymous):

A shortcut would be to simplify 2/4 to 1/2 and then multiply 1/2 by 3/3 to get 3/6 then add 3/6 and 5/6 to get 8/6, or 4/3.